OpenStudy (anonymous):

a+3-1x3+2-4+0x1=0 then a=

OpenStudy (anonymous):

is it -1x3

OpenStudy (anonymous):

or - 1x + 3

OpenStudy (anonymous):

a + 3 - 1 * 3 + 2 - 4 + 0 * 1 = 0 a + 3 - (1 * 3) + 2 - 4 + (0*1) = 0 a + 3 - 3 + 2 - 4 + 0 = 0 a + 3 - 3 + 2 - 4 = 0 a + 3 - 3 + 2 = 4 a + 3 - 3 = 2 a + 3 = 5 a = 2